• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh/JP

    Log entry time 19:21:45 on November29,2006

    Entry number 189488

    keyword=Hall access report for the Target problem

    JP and I took an access into the hall to fix the pressure of the hydrogen in loop3. With too much troubleshooting JP finally found a valve execcively hot which might have probably caused the leakage of the gas. The gas-pressure was continuously decreasing from more than 25 psi to 17,5 psi. JP opened the H2 gas cylinder there so as to supply more H2 into the system. Unfortunately the reserved gas there is running out and it needs to be replaced by some authorized personnel (safety reason). Now the pressure looks stable at 18.33 psi. When, in case, it goes below 16.5 psi TO needs to page Mike Seely and update him the situation. Otherwise keep it as it is until the experts pay proper attention.

    Below is the current target status.

    FIGURE 1